OpenStudy (shreehari499):

A step down transformer has 50 turns on secondary and 1000 turns on primary winding. If a transformer is connected to 220V 1A AC source, what is output current of the transformer?

OpenStudy (shreehari499):

@ganeshie8 ..:)

OpenStudy (anonymous):

\[\frac{ Ns }{ Np } = \frac{ Ip }{ Is }\]

OpenStudy (anonymous):

\[\frac{ 50 }{ 1000 } = \frac{ 1 }{ Is }\]

OpenStudy (anonymous):

Is = 20 A